The following is cost information for the Creamy Crisp Donut Company: Entrepreneur's potential earnings as a salaried worker = $50,000 Annual lease on building = $22,000 Annual revenue from operations = $380,000 Payments to workers = $120,000 Utilities (electricity, water, disposal) costs = $8,000 Value of entrepreneur's talent in the next best entrepreneurial activity = $80,000 Entrepreneur's forgone interest on personal funds used to finance the business = $6,000

20). Refer to the above data. Creamy Crisp's normal profit is: A. $160,000. B. $90,000. C. $80,000. D. $130,000.
Answer #1

Ans:

Explanation:

Explicit costs = Annual lease on building + Payments to workers + Utilities

                      = 22,000 + 120,000 + 8,000

                      = $150,000

Implicit costs = Entrepreneur's potential earnings as a salaried worker + Value of entrepreneur's talent in the next best entrepreneurial activity + Entrepreneur's forgone interest on personal funds used to finance the business

                      = 50,000 + 80,000 + 6,000

                       = $136,000

Normal profit = Annual revenue from operations - Explicit costs - Implicit costs

                      = 380,000 - 150,000 - 136,000

                      = $94,000